OpenStudy (liliakarina):

How many numbers of 4 digits can be formed out of {6,9}?

OpenStudy (anonymous):

choice of two numbers in each of 4 spots so by the "counting principle' you have \[2\times 2\times 2\times2=2^4=16\] choices

OpenStudy (amistre64):

its binomial in nature i think

OpenStudy (amistre64):

6666 6669 6696 6966 9666 6699 6969 9669 etc

OpenStudy (amistre64):

16 sounds good tho :)

OpenStudy (liliakarina):

So, 16 choices. Thanks

OpenStudy (anonymous):

nine down, seven to go...

OpenStudy (amistre64):

1+4+6+4+1 = 2+8+6 .... yeah, 16